OpenStudy (anonymous):

I need help in finding the range of the following function: arcsin(3x+1) For finding the domain I did the following... -1

Vocaloid (vocaloid):

yup, looks good to me! great job ~

OpenStudy (anonymous):

Great! Thanks @Vocaloid :) Ant ideas on how to find the range?

OpenStudy (anonymous):

Any*

Vocaloid (vocaloid):

Hm, that's a bit trickier, I suppose you could evaluate arcsin(-2/3) and arcsin(0), but I'm sure there's a better way, let me think about this one for a second

OpenStudy (anonymous):

Alrighty! :D Thanks so much!

OpenStudy (anonymous):

I think the range of arcsin x is always -π/2 ≤ y ≤ π/2

OpenStudy (anonymous):

Wait would it be pi/2 - 3pi/2 ?

jimthompson5910 (jim_thompson5910):

I think you meant to say this for the domain \[\Large -1 \le 3x{\color{red}+}1 \le 1\] and you solve for x from there

Vocaloid (vocaloid):

yes, peachpi is right, silly me!

OpenStudy (anonymous):

@jim_thompson5910 you caught my typo again >.<

OpenStudy (anonymous):

@peachpi thank you!

jimthompson5910 (jim_thompson5910):

but you have the correct result after you isolate x \[\Large -\frac{2}{3} \le x \le 0\] is the correct domain

OpenStudy (anonymous):

you're welcome

OpenStudy (anonymous):

@jim_thompson5910 okay so would be correct to find the range by just plugging in -2/3 and 0 ?

OpenStudy (anonymous):

*would it

jimthompson5910 (jim_thompson5910):

yep, plugging in x = -2/3 will give -pi/2 which is the lowest you can go in the range plugging in x = 0 will give pi/2 which is the highest you can go in the range